# Preference-based Reinforcement Learning beyond Pairwise Comparisons: Benefits of Multiple Options

### v1

@@ E.2 Proofs of Lemmas for Theorem 2
$\succeq\!\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\!\!\sum_{j=1}^{|S_{s}|-1}\!\!\sum_{k=j+1}^{|S_{s}|}\dot{\mu}\left(z_{sjk}^{\top}\bm{\theta}^{\star}\right)e^{-\left|z_{sjk}^{\top}\big(\widehat{\bm{\theta}}_{s}^{(j,k+1)}-\bm{\theta}^{\star}\big)\right|}z_{sjk}z_{sjk}^{\top}+\lambda\mathbf{I}_{d}$ (Lemma F.2)
$\succeq e^{-1}\!\!\!\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\!\!\sum_{j=1}^{|S_{s}|-1}\!\!\sum_{k=j+1}^{|S_{s}|}\dot{\mu}\left(z_{sjk}^{\top}\bm{\theta}^{\star}\right)z_{sjk}z_{sjk}^{\top}+\lambda\mathbf{I}_{d},$ (E.3)
where the last inequality holds because, for any $s\notin{\mathcal{T}}^{w}$, the following property is satisfied:
$\left|z_{sjk}^{\top}\big(\widehat{\bm{\theta}}_{s}^{(j,k+1)}-\bm{\theta}^{\star}\big)\right|$ $=\left|\big(\phi(x_{s},\sigma_{sj})-\phi(x_{s},\sigma_{sk})\big)^{\top}\left(\widehat{\bm{\theta}}_{s}^{(j,k+1)}-\bm{\theta}^{\star}\right)\right|$
$\leq\big\|\phi(x_{s},\sigma_{sj})-\phi(x_{s},\sigma_{sk})\big\|_{H_{s}^{-1}}\big\|\widehat{\bm{\theta}}_{s}^{(j,k+1)}-\bm{\theta}^{\star}\big\|_{H_{s}}$ (Hölder’s inequality)
@@
$\leq\frac{\beta_{t}(\delta)}{\beta_{T+1}(\delta)}$ (Corollary E.1)
$\leq 1.$ ($\beta_{t}(\delta)$ is non-decreasing)
For simplicity, we write $\mathbb{P}_{s}(a\succ a^{\prime})=\mathbb{P}(a\succ a^{\prime}|x_{s};\bm{\theta}^{\star})$. Let $P_{s,\{a,a^{\prime}\}}$ denote the Bernoulli distribution over the support $\{a,a^{\prime}\}$, where $a$ occurs with probability $\mu\big(\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime})\big)^{\top}\bm{\theta}^{\star}\big)$. Then, to further lower bound the right-hand side of Equation (E.3), we proceed as follows:
$\sum_{j=1}^{|S_{s}|-1}\!\!\sum_{k=j+1}^{|S_{s}|}\dot{\mu}\left(z_{sjk}^{\top}\bm{\theta}^{\star}\right)z_{sjk}z_{sjk}^{\top}$
$=\sum_{j=1}^{|S_{s}|-1}\!\!\sum_{k=j+1}^{|S_{s}|}\mu\left(z_{sjk}^{\top}\bm{\theta}^{\star}\right)\mu\left(z_{skj}^{\top}\bm{\theta}^{\star}\right)z_{sjk}z_{sjk}^{\top}$
$=\frac{1}{2}\sum_{a\in S_{s}}\sum_{a^{\prime}\in S_{s}}\mathbb{P}_{s}(a\succ a^{\prime})\mathbb{P}_{s}(a^{\prime}\succ a)\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime})\big)\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime})\big)^{\top}$
$\succeq\frac{1}{2}\sum_{a\in S_{s}}2\mathbb{P}_{s}(a\succ\bar{a}_{s})\mathbb{P}_{s}(\bar{a}_{s}\succ a)\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)^{\top}$ ($\bar{a}_{s}\in S_{s}$ by Eqn. (9))
$=\frac{1}{2}\sum_{a\in S_{s}}\mathbb{E}_{(a^{\prime},a^{\prime\prime})\sim P_{s,\{a,\bar{a}_{s}\}}^{\otimes 2}}\left[\big(\phi(x_{s},a^{\prime})-\phi(x_{s},a^{\prime\prime})\big)\big(\phi(x_{s},a^{\prime})-\phi(x_{s},a^{\prime\prime})\big)^{\top}\right],$ (E.4)
where $P_{s,\{a,\bar{a}_{s}\}}^{\otimes 2}=P_{s,\{a,\bar{a}_{s}\}}\times P_{s,\{a,\bar{a}_{s}\}}$ denotes the the product distribution over two independent samples from $P_{s,\{a,\bar{a}_{s}\}}$. Note that the Bernoulli distribution $P_{s,\{a,\bar{a}_{s}\}}$, where $a\in S_{s}$, is measurable with respect to the filtration $\mathcal{F}_{s-1}=\bm{\sigma}\left(S_{1},\sigma_{1},\dots,S_{s-1},\sigma_{s-1},S_{s}\right)$. Then, plugging Equation (E.4) into Equation (E.3), we get
$H_{t}$
$\succeq\frac{e^{-1}}{2}\!\!\!\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{a\in S_{s}}\mathbb{E}_{(a^{\prime},a^{\prime\prime})\sim P_{s,\{a,\bar{a}_{s}\}}^{\otimes 2}}\left[\big(\phi(x_{s},a^{\prime})-\phi(x_{s},a^{\prime\prime})\big)\big(\phi(x_{s},a^{\prime})-\phi(x_{s},a^{\prime\prime})\big)^{\top}\right]+\lambda\mathbf{I}_{d}$
$\succeq\frac{3e^{-1}}{10}\underbrace{\left(\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{a\in S_{s}}\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)^{\top}+\lambda\mathbf{I}_{d}\right)}_{=\Lambda_{t}}$ (covariance matrix concentration lemma (Corollary F.1))
$\succeq\frac{1}{10}\Lambda_{t},$
which conclude the proof of Lemma E.1. ∎
E.2.2 Proof of Lemma E.2
Proof of Lemma E.2.
For simplicity, let $\widetilde{{\mathcal{T}}}^{w}_{t}=\{s\in[t-1]\mid s\in{\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}\}$. Clearly, $\widetilde{{\mathcal{T}}}^{w}_{T+1}={\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}$. Recall that by the definition of $H_{t}$, we have
@@
$\succeq\kappa\sum_{s=1}^{t-1}\sum_{j=1}^{|S_{s}|-1}\!\!\sum_{k=j+1}^{|S_{s}|}z_{sjk}z_{sjk}^{\top}+\lambda\mathbf{I}_{d}$ ($\kappa=e^{-4B}/4$)
$\succeq\kappa\sum_{s=1}^{t-1}\sum_{a\in S_{s}}\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)^{\top}+\lambda\mathbf{I}_{d}$ ($\bar{a}_{s}\in S_{s}$)
$\succeq\kappa\,\underbrace{\left(\sum_{s\in\widetilde{{\mathcal{T}}}^{w}_{t}}\sum_{a\in S_{s}}\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)^{\top}+\lambda\mathbf{I}_{d}\right)}_{=:\Lambda^{w}_{t}}.$ (E.5)
Let $\tilde{a}_{t}=\mathop{\mathrm{argmax}}_{a\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|_{H_{t}^{-1}}$. Then, we get
$\sum_{t\in\widetilde{{\mathcal{T}}}^{w}_{T+1}}\max_{a\in\mathcal{A}}$ $\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|^{2}_{H_{t}^{-1}}$
$\leq\sum_{t\in\widetilde{{\mathcal{T}}}^{w}_{T+1}}\|\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\|^{2}_{H_{t}^{-1}}$
$\leq\sum_{t\in\widetilde{{\mathcal{T}}}^{w}_{T+1}}\sum_{a\in S_{t}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|^{2}_{H_{t}^{-1}}$ ($\tilde{a}_{t},\bar{a}_{t}\in S_{t}$ by Eqn. (9))
$\leq\frac{1}{\kappa}\sum_{t\in\widetilde{{\mathcal{T}}}^{w}_{T+1}}\sum_{a\in S_{t}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|^{2}_{\big(\Lambda^{w}_{t}\big)^{-1}}$ (Eqn. (E.5))
$\leq\frac{1}{\kappa}\sum_{t=1}^{T}\min\left\{1,\sum_{a\in S_{t}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|^{2}_{\big(\Lambda^{w}_{t}\big)^{-1}}\right\}$ ($t\neq{\mathcal{T}}_{0}$ and $\widetilde{{\mathcal{T}}}^{w}_{T+1}\subseteq[T]$)
$\leq\frac{2}{\kappa}d\log\left(1+\frac{2KT}{d\lambda}\right).$ (Lemma D.3)
On the other hand, for $t\in\widetilde{{\mathcal{T}}}^{w}_{T+1}={\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}$, we know that
$\sum_{t\in\widetilde{{\mathcal{T}}}^{w}_{T+1}}\max_{a\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|^{2}_{H_{t}^{-1}}\geq\frac{|\widetilde{{\mathcal{T}}}^{w}_{T+1}|}{\beta_{T+1}(\delta)^{2}}.$
By combining the two results above, we get
$|\widetilde{{\mathcal{T}}}^{w}_{T+1}|\leq\frac{2}{\kappa}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{2KT}{d\lambda}\right),$
@@
Appendix F Technical Lemmas
Lemma F.1 (Proposition B.5 of Lee and Oh 43).
@@ Appendix F Technical Lemmas
Lemma F.2 (Lemma 9 of Abeille et al. 2).
@@
Lemma F.3 (Concentration of covariances, Lemma 39 of Zanette et al. 92).
Let $\mu_{i}$ be the conditional distribution of $\phi\in\mathbb{R}^{d}$ given the sampled $\phi_{1},\dots,\phi_{i-1}$. Assume $\|\phi\|_{2}\leq 1$. Define $\Sigma=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\phi\sim\mu_{i}}\phi\phi^{\top}$. If $\lambda=\Omega(d\log(n/\delta))$, then, with probability at least $1-\delta$, for any $n\geq 1$, we have
$\frac{1}{3}\left(n\Sigma+\lambda\mathbf{I}_{d}\right)\preceq\sum_{i=1}^{n}\phi_{i}\phi_{i}^{\top}+\lambda\mathbf{I}_{d}\preceq\frac{5}{3}\left(n\Sigma+\lambda\mathbf{I}_{d}\right).1$
We now extend the previous sequence $\phi_{1},\dots,\phi_{i-1}$ to a general setting where conditioning is performed with respect to an arbitrary filtration $\mathcal{F}_{i-1}$. For instance, $\mathcal{F}_{i-1}$ may be the $\sigma$-algebra generated by previous samples $\tilde{\phi}_{1},\dots,\tilde{\phi}_{i-1}$, where each $\tilde{\phi}_{j}$ is drawn from a potentially different distribution $\tilde{\mu}_{j}$. This generalization is valid because a martingale difference sequence can be defined with respect to any filtration provided that the $\sigma$-algebras satisfy the usual properties (e.g., nestedness and making $\{\phi_{i}\}$ adapted).
Corollary F.1 (Generalized version of covariance concentration).
Let $\mu_{i}$ denote the conditional distribution of $\phi\in\mathbb{R}^{d}$ conditioned on the filtration $\mathcal{F}_{i-1}$. Assume $\|\phi\|_{2}\leq 1$. Define $\Sigma=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\phi\sim\mu_{i}}\phi\phi^{\top}$. If $\lambda=\Omega(d\log(n/\delta))$, then with probability at least $1-\delta$, for any $n\geq 1$, we have
$\frac{1}{3}\left(n\Sigma+\lambda\mathbf{I}_{d}\right)\preceq\sum_{i=1}^{n}\phi_{i}\phi_{i}^{\top}+\lambda\mathbf{I}_{d}\preceq\frac{5}{3}\left(n\Sigma+\lambda\mathbf{I}_{d}\right).$
We also provide a concentration lemma for the more general case of positive semi-definite (PSD) random matrices.
Lemma F.4 (Concentration of PSD matrices).
Let $\mu_{i}$ denote the conditional distribution of a positive semi-definite $M\in\mathbb{R}^{d\times d}$ conditioned on the filtration $\mathcal{F}_{i-1}$. Assume $\lambda_{\max}(M)\leq 1$. Define $\widebar{M}:=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{M\sim\mu_{i}}M$. If $\lambda=\Omega(d\log(n/\delta))$, then with probability at least $1-\delta$, for any $n\geq 1$, we have
@@
Proof of Lemma F.4.
The overall structure of the proof closely follows that of Lemma 39 in Zanette et al. 92. For completeness, we provide the full proof below.
@@
Since $M$ is a positive semi-definite matrix, the random variable $x^{\top}Mx$ is non-negative with the maximum value $x^{\top}Mx\leq\lambda_{\max}(M)\|x\|_{2}^{2}\leq 1$. Thus, the conditional variance is at most $x^{\top}\widebar{M}_{i}x$ because
@@
Then, by Lemma F.5, with probability at least $1-\delta$, for some constant $c$, we have
$\left|\frac{1}{n}\sum_{i=1}^{n}\left(x^{\top}M_{i}x-x^{\top}\widebar{M}_{i}x\right)\right|=\left|\frac{1}{n}\sum_{i=1}^{n}x^{\top}M_{i}x-x^{\top}\widebar{M}x\right|\leq c\left(\sqrt{2\frac{x^{\top}\widebar{M}x}{n}\log(2/\delta)}+\frac{\log(2/\delta)}{3n}\right).$
@@
$c\left(\sqrt{2\frac{x^{\top}\widebar{M}x}{n}\log(2/\delta)}+\frac{\log(2/\delta)}{3n}\right)\leq\frac{1}{2}\left(x^{\top}\widebar{M}x+\frac{\lambda}{n}\right).$ (F.1)
@@
$\sqrt{2\frac{\log(2/\delta)}{n}}\leq c^{\prime}\sqrt{\frac{\lambda}{n}}\quad$ $\longleftrightarrow\quad\Omega(\log(1/\delta))\leq\lambda$
@@
$\sqrt{2\frac{x^{\top}\widebar{M}x}{n}\log(2/\delta)}\leq c^{\prime\prime\prime}\left(\frac{\lambda}{n}\right)\quad$ $\longleftrightarrow\quad\Omega(\log(1/\delta))\leq\lambda$
@@
Therefore, Equation (F.1) is satisfied. Since $\|x\|_{2}\leq 1$, this implies
$\left|x^{\top}\left(\frac{1}{n}\sum_{i=1}^{n}M_{i}-\widebar{M}\right)x\right|\leq\frac{1}{2}x^{\top}\left(\widebar{M}+\frac{\lambda}{n}\mathbf{I}_{d}\right)x.$ (F.2)
We denote the boundary of the unit ball by $\partial\mathcal{B}=\left\{\|x\|_{2}=1\right\}$. Then, for any $x\in\partial\mathcal{B}$, we know there exists a $x^{\prime}$ in the $\epsilon$-covering such that $\|x-x^{\prime}\|_{2}\leq\epsilon$. Let $\mathcal{N}_{\epsilon}$ be the $\epsilon$-covering number of $\partial\mathcal{B}$. Then, by the covering number of Euclidean ball lemma (Lemma F.6), we get
$\mathcal{N}_{\epsilon}\leq\left(\frac{3}{\epsilon}\right)^{d}.$ (F.3)
@@
$\leq\frac{1}{2}(x^{\prime})^{\top}\left(\widebar{M}+\frac{\lambda}{n}\mathbf{I}_{d}\right)x^{\prime}+4\epsilon$ (Eqn. (F.2))
@@
where $\lambda=\Omega\left(\log\left(\frac{2n\mathcal{N}_{\epsilon}}{\delta}\right)\right)$. By substituting $\delta\leftarrow\delta/(n\mathcal{N}_{\epsilon}+1)$ and combining this with Equation (F.3), we obtain:
@@
Lemma F.5 (Bernstein for martingales, Theorem 1 of Beygelzimer et al. 10 and Lemma 45 of Zanette et al. 92).
@@
Lemma F.6 (Covering number of Euclidean ball).
@@
Appendix G Proof of Theorem 3
G.1 Main Proof of Theorem 3

### v2

@@ D.2 Proofs of Lemmas for Theorem 2
$\succeq\!\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\!\!\sum_{j=1}^{|S_{s}|-1}\!\!\sum_{k=j+1}^{|S_{s}|}\dot{\mu}\left(z_{sjk}^{\top}\widehat{\bm{\theta}}_{s}\right)e^{-\left|z_{sjk}^{\top}\big(\widehat{\bm{\theta}}_{s}^{(j,k+1)}-\widehat{\bm{\theta}}_{s}\big)\right|}z_{sjk}z_{sjk}^{\top}+\lambda\mathbf{I}_{d}$ (Lemma E.2)
$\succeq e^{-2}\!\!\!\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\!\!\sum_{j=1}^{|S_{s}|-1}\!\!\sum_{k=j+1}^{|S_{s}|}\dot{\mu}\left(z_{sjk}^{\top}\widehat{\bm{\theta}}_{s}\right)z_{sjk}z_{sjk}^{\top}+\lambda\mathbf{I}_{d}$
$\succeq\frac{e^{-2}}{2}\Lambda_{t},$
where the second-to-last inequality holds because, for any $s\notin{\mathcal{T}}^{w}$, the following property is satisfied:
$\left|z_{sjk}^{\top}\big(\widehat{\bm{\theta}}_{s}^{(j,k+1)}-\widehat{\bm{\theta}}_{s}\big)\right|$ $=\left|\big(\phi(x_{s},\sigma_{sj})-\phi(x_{s},\sigma_{sk})\big)^{\top}\left(\widehat{\bm{\theta}}_{s}^{(j,k+1)}-\widehat{\bm{\theta}}_{s}\right)\right|$
$\leq\big\|\phi(x_{s},\sigma_{sj})-\phi(x_{s},\sigma_{sk})\big\|_{H_{s}^{-1}}\left(\big\|\widehat{\bm{\theta}}_{s}^{(j,k+1)}-\bm{\theta}^{\star}\big\|_{H_{s}}+\big\|\widehat{\bm{\theta}}_{s}-\bm{\theta}^{\star}\big\|_{H_{s}}\right)$ (Hölder’s inequality)
@@
$\leq\frac{2\beta_{T+1}(\delta)}{\beta_{T+1}(\delta)}$ (Corollary D.1 and $\beta_{t}(\delta)$ is non-decreasing)
$\leq 2.$
This concludes the proof of Lemma D.1. ∎
D.2.2 Proof of Lemma D.2
Proof of Lemma D.2.
The proof follows the same structure as the analysis of Lemma C.4, with the main difference that we use the Hessian corresponding to the RB loss. Let
$(\tilde{a}_{t},\bar{a}_{t})=\mathop{\mathrm{argmax}}_{a,a^{\prime}\in\mathcal{A}}\dot{\mu}\left(\big(\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|_{H_{t}^{-1}}^{2}.$
By the assortment selection rule in (11), the two actions $\tilde{a}_{t},\bar{a}_{t}$ are always included in $S_{t}$. It is because, for any pair $\{a,a^{\prime}\}$, we have
$\dot{\mu}\left(\big(\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|^{2}_{H_{t}^{-1}}=2f_{t}(\{a,a^{\prime}\}),$
which directly implies that $\tilde{a}_{t},\bar{a}_{t}\in S_{t}$.
For simplicity, we denote $z_{skk^{\prime}}=\phi(x_{s},\sigma_{sk})-\phi(x_{s},\sigma_{sk^{\prime}})$. Recall that by the definition of $H_{t}$,
@@
$\succeq\frac{e^{-4B}}{4}\sum_{s=1}^{t-1}\sum_{j=1}^{|S_{s}|-1}\!\!\sum_{k=j+1}^{|S_{s}|}z_{sjk}z_{sjk}^{\top}+\lambda\mathbf{I}_{d}$
$\succeq\frac{e^{-4B}}{4}\sum_{s=1}^{t-1}\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)^{\top}+\lambda\mathbf{I}_{d}$ ( $\tilde{a}_{s},\bar{a}_{s}\in S_{s}$)
$\succeq\frac{e^{-4B}}{4}\Bigg(\sum_{s\in{\mathcal{T}}^{w}_{t}}\dot{\mu}\!\left(\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)^{\top}\widehat{\bm{\theta}}_{s}\right)\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)^{\top}$
$\quad\quad\quad\quad+\lambda\mathbf{I}_{d}\Bigg),$ (D.5)
where in the last inequality, we use the fact that $\dot{\mu}(\cdot)\leq\frac{1}{4}$ and let ${\mathcal{T}}^{w}_{t}={\mathcal{T}}^{w}\setminus\{t,\dots,T\}$. Define
$\Lambda^{w}_{t}:=$ $\sum_{s\in{\mathcal{T}}^{w}_{t}}\dot{\mu}\!\left(\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)^{\top}\widehat{\bm{\theta}}_{s}\right)\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)^{\top}$
$+\lambda\mathbf{I}_{d}.$
Furthermore, by setting $\lambda\geq 1$, for any $t$, the following holds
$\dot{\mu}\left(\big(\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\|\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\|^{2}_{\big(\Lambda^{w}_{t}\big)^{-1}}$ $\leq\frac{1}{4}\cdot\frac{2}{\lambda}$ ($\dot{\mu}(\cdot)\leq\frac{1}{4}$)
$\leq 1.$ (D.6)
Then, we obtain
$\sum_{t\in{\mathcal{T}}^{w}}$ $\max_{a,a^{\prime}\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|^{2}_{H_{t}^{-1}}$
$=\sum_{t\in{\mathcal{T}}^{w}}\max_{a,a^{\prime}\in\mathcal{A}}\frac{\dot{\mu}\left(\big(\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)}{\dot{\mu}\left(\big(\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)}\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|^{2}_{H_{t}^{-1}}$
$\leq 4e^{2B}\sum_{t\in{\mathcal{T}}^{w}}\max_{a,a^{\prime}\in\mathcal{A}}\dot{\mu}\left(\big(\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|^{2}_{H_{t}^{-1}}$ (Assumption 1)
$=4e^{2B}\sum_{t\in{\mathcal{T}}^{w}}\dot{\mu}\left(\big(\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\|\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\|^{2}_{H_{t}^{-1}}$
$\leq 16e^{6B}\sum_{t\in{\mathcal{T}}^{w}}\dot{\mu}\left(\big(\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\|\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\|^{2}_{\big(\Lambda^{w}_{t}\big)^{-1}}$ (Eqn. (D.5))
$=16e^{6B}\sum_{t\in{\mathcal{T}}^{w}}\min\left\{1,\dot{\mu}\left(\big(\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\|\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\|^{2}_{\big(\Lambda^{w}_{t}\big)^{-1}}\right\}$ (Eqn. (D.6) )
$\leq 32e^{6B}d\log\left(1+\frac{2T}{d\lambda}\right).$ (Lemma E.6)
On the other hand, for $t\in{\mathcal{T}}^{w}$, we know that
$\sum_{t\in{\mathcal{T}}^{w}}\max_{a,a^{\prime}\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|^{2}_{H_{t}^{-1}}\geq\frac{|{\mathcal{T}}^{w}|}{\beta_{T+1}(\delta)^{2}}.$
By combining the two results above, and setting $\kappa=e^{-6}$, we derive that
$|{\mathcal{T}}^{w}|\leq\frac{32}{\kappa}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{2T}{d\lambda}\right),$
@@
Appendix E Technical Lemmas
Lemma E.1 (Proposition B.5 of lee2025improved).
@@ Appendix E Technical Lemmas
Lemma E.2 (Lemma 9 of abeille2021instance).
@@
We also provide a concentration lemma for positive semi-definite (PSD) random matrices.
Lemma E.3 (Concentration of PSD matrices).
Let $\mu_{i}$ denote the conditional distribution of a positive semi-definite $M\in\mathbb{R}^{d\times d}$ conditioned on the filtration $\mathcal{F}_{i-1}$. Assume $\lambda_{\max}(M)\leq 1$. Define $\widebar{M}:=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{M\sim\mu_{i}}M$. If $\lambda=\Omega(d\log(n/\delta))$, then with probability at least $1-\delta$, for any $n\geq 1$,
@@
Proof of Lemma E.3.
The overall structure of the proof closely follows that of Lemma 39 in zanette2021cautiously. For completeness, we provide the full proof below.
@@
Since $M$ is a positive semi-definite matrix, the random variable $x^{\top}Mx$ is non-negative, and it satisfies $x^{\top}Mx\leq\lambda_{\max}(M)\|x\|_{2}^{2}\leq 1$. Thus, the conditional variance is at most $x^{\top}\widebar{M}_{i}x$ because
@@
Applying Lemma E.4 with the filtration $\mathcal{F}_{i}$, we obtain that, with probability at least $1-\delta$, there exists a universal constant $c$ such that
$\left|\frac{1}{n}\sum_{i=1}^{n}\left(x^{\top}M_{i}x-x^{\top}\widebar{M}_{i}x\right)\right|=\left|\frac{1}{n}\sum_{i=1}^{n}x^{\top}M_{i}x-x^{\top}\widebar{M}x\right|\leq c\left(\sqrt{\frac{2x^{\top}\widebar{M}x}{n}\log(2/\delta)}+\frac{\log(2/\delta)}{3n}\right).$
@@
$c\left(\sqrt{\frac{2x^{\top}\widebar{M}x}{n}\log(2/\delta)}+\frac{\log(2/\delta)}{3n}\right)\leq\frac{1}{2}\left(x^{\top}\widebar{M}x+\frac{\lambda}{n}\right).$ (E.1)
@@
$\sqrt{\frac{2\log(2/\delta)}{n}}\leq c^{\prime}\sqrt{\frac{\lambda}{n}}\quad$ $\longleftrightarrow\quad\Omega(\log(1/\delta))\leq\lambda$
@@
$\sqrt{\frac{2x^{\top}\widebar{M}x}{n}\log(2/\delta)}\leq c^{\prime\prime\prime}\left(\frac{\lambda}{n}\right)\quad$ $\longleftrightarrow\quad\Omega(\log(1/\delta))\leq\lambda$
@@
Therefore, Equation (E.1) is satisfied. Since $\|x\|_{2}\leq 1$, this implies
$\left|x^{\top}\left(\frac{1}{n}\sum_{i=1}^{n}M_{i}-\widebar{M}\right)x\right|\leq\frac{1}{2}x^{\top}\left(\widebar{M}+\frac{\lambda}{n}\mathbf{I}_{d}\right)x.$ (E.2)
We denote the boundary of the unit ball by $\partial\mathcal{B}=\left\{\|x\|_{2}=1\right\}$. Then, for any $x\in\partial\mathcal{B}$, we know there exists a $x^{\prime}$ in the $\epsilon$-covering such that $\|x-x^{\prime}\|_{2}\leq\epsilon$. Let $\mathcal{N}_{\epsilon}$ be the $\epsilon$-covering number of $\partial\mathcal{B}$. Then, by the covering number of Euclidean ball lemma (Lemma E.5), we get
$\mathcal{N}_{\epsilon}\leq\left(\frac{3}{\epsilon}\right)^{d}.$ (E.3)
@@
$\leq\frac{1}{2}(x^{\prime})^{\top}\left(\widebar{M}+\frac{\lambda}{n}\mathbf{I}_{d}\right)x^{\prime}+4\epsilon$ (Eqn. (E.2))
@@
where $\lambda=\Omega\left(\log\left(\frac{2n\mathcal{N}_{\epsilon}}{\delta}\right)\right)$. By substituting $\delta\leftarrow\delta/(n\mathcal{N}_{\epsilon}+1)$ and combining this with Equation (E.3), we obtain:
@@
Lemma E.4 (Bernstein for martingales, Theorem 1 of beygelzimer2011contextual and Lemma 45 of zanette2021cautiously).
@@
Lemma E.5 (Covering number of Euclidean ball).
@@
Lemma E.6 (Elliptical potential lemma abbasi2011improved).
Let $\{z_{t}\}_{t\geq 1}$ be a bounded sequence in $\mathbb{R}^{d}$ satisfying $\max_{t\geq 1}\|z_{t}\|_{2}\leq X$. For any $t\geq 1$, we define $\Lambda_{t}:=\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}+\lambda\mathbf{I}_{d}$ with $\lambda>0$. Then, we have
$\sum_{t=1}^{T}\min\left\{1,\|z_{t}\|_{\Lambda_{t}^{-1}}^{2}\right\}\leq 2d\log\left(1+\frac{X^{2}T}{d\lambda}\right).$
Appendix F Proof of Theorem 3
F.1 Main Proof of Theorem 3